# Revisiting generic mean curvature flow in $\mathbb{R}^{3}$

### v2

@@ 1. Introduction
Here, $\mathbf{H}_{M(t)}(\mathbf{x})$ denotes the mean curvature vector of $M(t)$ at $\mathbf{x}$. By a simple application of the maximum principle, the flow $M(t)$ becomes singular in finite time whenever $M(0)\subset\mathbb{R}^{n+1}$ is closed and embedded. There are several weak notions of mean curvature flow that allow to flow through singularities. In this paper, we use the notion of unit-regular Brakke flows [Bra75], but see also Remark 1.
A well-known conjecture of Huisken (see [Ilm03, #8]) posited that generic mean curvarture flows in $\mathbb{R}^{3}$ encounter only spherical and cylindrical singularities. The authors proved this conjecture under the hypothesis that all singularities arise with multiplicity one, by developing a one-sided perturbation and genus-drop mechanism [CCMS24a, CCS23]. This assumption that singularities always have multiplicity one (Ilmanen’s multiplicity one conjecture [Ilm03, #2]) was recently proven to always hold in a remarkable work of Bamler–Kleiner [BK23]. As such, when combined with our previous work, this resolved Huisken’s conjecture.
@@
[…]
Here are some well-known implications of ()($\dagger$)( † ):
@@
    The work of Hershkovits–White [HW20], combined with the resolution of the mean convex neighborhood conjecture by Choi–Haslhofer–Hershkovits [CHH22], yields that the level-set flow of $M$ is non-fattening and that the innermost and outermost flows starting at $M$ agree, with their spacetime support equal to the level-set flow of $M$. Moreover, $\mathcal{M}$ is the unique cyclic unit-regular Brakke flow starting at $M$ by [CCMS24a, Appendix G]. Therefore $\mathcal{M}$, the innermost, and the outermost flows starting at $M$ all coincide and are almost-regular in the sense of Bamler–Kleiner [BK23, Definition 2.6] (cf. [BK23, Lemma 7.8]).
@@
    Consider any sequence $M_{i}\to M$, say in $C^{1}$, and a corresponding sequence $\mathcal{M}_{i}$ of cyclic unit-regular Brakke flows starting at $M_{i}$. Since the class of cyclic unit-regular Brakke flows is closed under weak convergence of Brakke flows, it follows from (a) above that $\mathcal{M}_{i}\rightharpoonup\mathcal{M}$. Furthermore, it follows from the classification of low entropy shrinkers [BW17] (or, alternatively, the proof of the mean convex neighborhood conjecture–see [SS20, Propsition 2.3]) that, for sufficiently large $i$, $\mathcal{M}_{i}$ also only has only multiplicity-one spherical and cylindrical singularities. In particular, the set of $M$ with corresponding $\mathcal{M}$ satisfying ($\dagger$1) is both dense (per Theorem 1.1) and open.
@@ 1.2. The genus-drop approach
The first step is to now study the rescaled mean curvature flow. By developing a PDE classification tool for solutions with fixed-sign speed (positive or negative) of the rescaled mean curvature flow on one side of non-spherical and non-cylindrical shrinkers, we show that the blow-up limits of the nearby $\mathcal{M}_{s}$ move monotonically in the sense of rescaled flows (i.e. in a shrinker mean convex sense in the original scale). This built on previous work of the authors in [CM22], inspired by [ADS19, ADS20]. One then deduces from this, as in [Whi00], that the $(\mathcal{M}_{s})_{s\neq 0}$ only develop multiplicity-one spherical and cylindrical singularities and have genus zero near $(\mathbf{x}_{0},t_{0})$.
@@
This genus-drop approach to resolve Huisken’s conjecture is significantly longer than the density-drop approach presented in the current paper, but it also gives a refined local picture. In particular, together with the Bamler–Kleiner multiplicity-one theorem and their compactness theorem for bounded almost regular flows ([BK23, Theorem 1.7]), inner/outermost flows experience strict genus drop after every nongeneric singular point (thus bounding the number of nongeneric singular points by the initial genus) see, e.g., the proof of [BK23, Theorem 1.9 (c)].
@@ 1.4. Acknowledgements
OC was supported by Terman Fellowship and an NSF grant (DMS-2304432). KC was supported by the KIAS Individual Grant MG078902. CM was supported by grant NSF DMS 2403728.
@@ 2.1. Spacetime
Furthermore we denote with
$d_{p}(X,Y)=d_{p}\big{(}(\mathbf{x},t),(\mathbf{y},s)\big{)}:=\sqrt{\|\mathbf{x}-\mathbf{y}\|^{2}+|t-s|}$
the parabolic distance on spacetime.
2.2. $F$-functional and density
@@
which is nondecreasing in $r$ by Huisken’s well-known monotonicity formula. In particular,
@@ 2.3. Shrinkers
$\mathcal{S}$ :=$:=\{\Sigma\subset\mathbb{R}^{3}\text{ embedded shrinker with }\lambda(\Sigma)<\infty\},$
@@ 2.4. Brakke flows
    For a bounded interval $[t_{1},t_{2}]\subset I$ and any compact set $K\subset\mathbb{R}^{3}$,
[…]
@@
    If $[t_{1},t_{2}]\subset I$, $f\in C^{1}_{c}(\mathbb{R}^{3}\times[t_{1},t_{2}])$, and $f\geq 0$, then Brakke’s inequality holds:
[…]
@@
In this weaker setting setting, (2.1), (2.2) extend to varifolds by looking at their induced Radon measures, and the monotonicity underlying (2.3), (2.4) extends to Brakke flows in $\mathbb{R}^{3}$ with bounded area ratios (cf. [Ilm95, Lemma 7]).
Furthermore, we denote with $\operatorname{ParDil}_{\lambda}$ the standard parabolic dilation of a Brakke flow around the spacetime origin $(\mathbf{0},0)$ by a factor $\lambda>0$.
@@ 2.5. $F$-stationary varifolds
The shrinker condition is also equivalent to the following (see [CM12a, §3]):
@@
A varifold $V$ in $\mathbb{R}^{3}$ is called $F$-stationary if is stationary with respect to the conformally flat metric $e^{-|\cdot|^{2}/4}g_{\mathbb{R}^{3}}$ on $\mathbb{R}^{3}$.
@@
~$\tilde{\mathcal{M}}_{i}:=\operatorname{ParDil}_{\lambda_{i}}(\mathcal{M}-X)\rightharpoonup\tilde{\mathcal{M}},$
@@
for some $F$-stationary integral varifold $V$ in $\mathbb{R}^{3}$.
@@ 3.1. Shrinker compactness
$\operatorname{genus}(\Sigma_{i})\leq g,\;\lambda(\Sigma_{i})\leq\Lambda,$
@@
$\Sigma_{i}\to\Sigma\text{ in }C^{\infty}_{\textnormal{loc}},\;F(\Sigma_{i})\to F(\Sigma).$
@@ 3.4. Frankel property for $F$-stationary varifolds
then $V=k\Sigma$ for some $k\in\mathbb{N}$ [^3] Here and throughout, we implicitly identify $\Sigma$ with its induced integral varifold.
@@ 3.6. First nongeneric singular time
We recall the following notion from [CCMS24a]: [^4] In the reference, $T_{\textnormal{bad}}(\mathcal{M})$ is equivalently defined as the supremum of times $T$ such that all tangent flows at $X=(\mathbf{x},t)\in\operatorname{sing}\mathcal{M}$, $t<T$, are for negative times concident with some $\mathcal{M}_{\Sigma}$, $\Sigma\in\mathcal{S}^{\textnormal{gen}}$.
@@ 3.7. Ilmanen and Bamler–Kleiner tangent flow structure theorem
The work of Ilmanen [Ilm95] and Bamler–Kleiner [BK23] gives us a precise description of tangent flows at the first nongeneric time time:
@@
Let $\mathcal{M}$ be a unit-regular integral Brakke flow with $\mathcal{M}(0)=\mathcal{H}^{2}\lfloor M$, where $M\subset\mathbb{R}^{3}$ is a closed embedded surface. If $T_{\textnormal{bad}}(\mathcal{M})<\infty$ and $X$ is a first nongeneric point for $\mathcal{M}$ (see Corollary 3.8), then every tangent flow to $\mathcal{M}$ at $X$ coincides, for $t<0$, with $\mathcal{M}_{\Sigma}$ for some $\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}$.
@@
The weaker result, with $\mathcal{M}_{\Sigma}$ replaced by $k\mathcal{M}_{\Sigma}$, $k\in\mathbb{N}$, and $\Sigma\in\mathcal{S}$, was proven in [Ilm95] for the first singular time. Modifications of the proof to the first nongeneric singular time were give in the proof of [CCMS24a, Proposition 11.3].
The fact that $k=1$ and $\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}$ follows from Bamler–Kleiner’s recent resolution of the multiplicity-one conjecture. In particular, it follows from [BK23, Theorem 1.2] provided we can ensure $\mathcal{M}$ is what Bamler–Kleiner call “almost-regular” on $[0,T_{\textnormal{bad}}(\mathcal{M}))$. This is justified in [BK23, Lemma 7.8]. ∎
@@
Proof.
@@
Using the compactness of cyclic unit-regular integral Brakke flows together with (c) and (4.4), we can pass (c) to the limit:
@@
The upper semicontinuity of densities together with (4.1), (4.6), (4.7), (4.8), implies
@@
In particular, $\mathcal{M}$ is a shrinking flow with spacetime center $X$ due to the monotonicity formula. If $V$ is the $F$-stationary varifold corresponding to some tangent flow to $\mathcal{M}$ at $\infty$- ∞, then
@@
It follows from Proposition 3.5, (4.10), and (4.11), that $V=\Sigma$. This forces Proposition 3.6 into a contradiction with (4.5). This completes our proof. ∎
@@
Denote $M_{s}:=\partial K_{s}$ and let $\mathcal{M}_{s}$ be any cyclic unit-regular integral Brakke flow with $\mathcal{M}_{s}(0)=\mathcal{H}^{2}\lfloor M_{s}$ (see [Ilm94]).
@@
Claim.
Fix $s_{0}\in(-1,1)$ and $X_{0}$ a first nongeneric point for $\mathcal{M}_{s_{0}}$ (recall Corollary 3.8). If $s_{0}\neq s_{i}\to s_{0}$, and $X_{i}\to X_{0}$ are first nongeneric points for $\mathcal{M}_{s_{i}}$, then
@@
for sufficiently large $i$. Translate so that $X_{0}=(\mathbf{0},0)$ and set $\lambda_{i}:=d_{p}(X_{i},(\mathbf{0},0))>0$; positivity follows from the avoidance principle [Whi95].
@@
(4.15) ~$\tilde{\mathcal{M}}_{s_{0}}^{i}:=\operatorname{ParDil}_{\lambda_{i}}\mathcal{M}_{s_{0}}\rightharpoonup\tilde{\mathcal{M}}_{s_{0}}=\mathcal{M}_{\Sigma}\text{for some }\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}.$
@@
By the nonfattening of the flow before $X_{0}$ by Hershkovits–White and Choi–Haslhofer–Hershkovits [HW20, CHH22], the regularity for a.e. $t$ of flows with only generic singularities by Colding–Minicozzi [CM16], and White’s genus monotonicity [Whi95],
@@
~$\tilde{\mathcal{M}}^{i}:=\operatorname{ParDil}_{\lambda_{i}}\mathcal{M}_{s_{i}}\rightharpoonup\tilde{\mathcal{M}},$
@@
$=\lim_{i}\Theta_{\mathcal{M}_{i}}(\lambda_{i}Y,\lambda_{i}r)$
$\leq\lim_{i}\Theta_{\mathcal{M}_{i}}(\lambda_{i}Y,\rho)$
@@
where $\Omega\subset\mathbb{R}^{3}$ is a component of $\mathbb{R}^{3}\setminus\Sigma$. This follows from (4.12) combined with the avoidance principle and the locally smooth convergence ~$\tilde{\mathcal{M}}^{i}_{s_{0}}\rightharpoonup\mathcal{M}_{\Sigma}$ on $\mathfrak{t}^{-1}((-\infty,0))$ from (4.15) and Brakke’s regularity theorem [Bra75], the latter of which applies since $\Sigma$ is smooth.
Altogether (4.16), (4.17), (4.18), (4.19), and (4.20) contradict our choice of $\eta_{0}$ coming from Proposition 4.1. This completes our proof of the claim. ∎
@@
$\mathcal{S}_{\ell}:=\bigcup_{s\in[-1,1]}$$\{X\text{ is a first nongeneric singular point for }\mathcal{M}_{s}$
@@
The claim implies that $\mathcal{S}_{\ell}$ is discrete. Therefore, $\cup_{\ell}\mathcal{S}_{\ell}$ is at most countable.

### v3

@@ 1. Introduction
Here, $\mathbf{H}_{M(t)}(\mathbf{x})$ denotes the mean curvature vector of $M(t)$ at $\mathbf{x}$. By a simple application of the maximum principle, the flow $M(t)$ becomes singular in finite time whenever $M(0)\subset\mathbb{R}^{n+1}$ is closed and embedded. There are several weak notions of mean curvature flow that allow to flow through singularities. In this paper, we use the notion of cyclic, unit-regular, integral Brakke flows. See Section 2 for definitions and references.
A well-known conjecture of Huisken (see [Ilm03, #8]) posited that generic mean curvarture flows in $\mathbb{R}^{3}$ encounter only spherical and cylindrical singularities. Groundbreaking progress toward Huisken’s conjecture was made in the work of Colding–Minicozzi [CM12a], who proved that spheres and cylinders are, in a sense, the only linearly stable singularity models for mean curvature flow (in all dimensions). They also showed how to construct “broken” mean curvature flows that avoid closed, multiplicity-one, non-spherical singular points. Subsequently, the present authors developed a one-sided perturbation and genus-drop mechanism [CCMS24a, CCS23] which proves Huisken’s conjecture under the hypothesis that all singularities arise with multiplicity one. This assumption that singularities always have multiplicity one (Ilmanen’s multiplicity one conjecture [Ilm03, #2]) was recently proven to always hold in a remarkable work of Bamler–Kleiner [BK23]. As such, when combined with our previous work, this resolved Huisken’s conjecture.
@@
($\dagger$)  $\mathcal{M}\text{ only has multiplicity-one spherical and cylindrical singularities}.$
Here are some well-known implications of $(\dagger)$:
@@
    The work of Hershkovits–White [HW20], combined with the resolution of the mean convex neighborhood conjecture by Choi–Haslhofer–Hershkovits [CHH22], yields that the level-set flow of $M$ is non-fattening and that the innermost and outermost flows starting at $M$ agree, with their spacetime support equal to the level-set flow of $M$. Moreover, $\mathcal{M}$ is the unique cyclic unit-regular integral Brakke flow starting at $M$ by [CCMS24a, Appendix G]. Therefore $\mathcal{M}$, the innermost, and the outermost flows starting at $M$ all coincide and are almost-regular in the sense of Bamler–Kleiner [BK23, Definition 2.6] (cf. [BK23, Lemma 7.8]).
@@
    Consider any sequence $M_{i}\to M$, say in $C^{1}$, and a corresponding sequence $\mathcal{M}_{i}$ of cyclic unit-regular integral Brakke flows starting at $M_{i}$. Since the class of cyclic unit-regular integral Brakke flows is closed under weak convergence of Brakke flows, it follows from (a) above that $\mathcal{M}_{i}\rightharpoonup\mathcal{M}$. Furthermore, it follows from the classification of low entropy shrinkers [BW17] (or, alternatively, the proof of the mean convex neighborhood conjecture–see [SS20, Propsition 2.3]) that, for sufficiently large $i$, $\mathcal{M}_{i}$ also only has only multiplicity-one spherical and cylindrical singularities. In particular, the set of $M$ with corresponding $\mathcal{M}$ satisfying ($\dagger$1) is both dense (per Theorem 1.1) and open.
@@ 1.2. The genus-drop approach
The first step is to now study the rescaled mean curvature flows of these blow ups. By developing a PDE classification tool for solutions with fixed-sign speed (positive or negative) of the rescaled mean curvature flow on one side of non-spherical and non-cylindrical shrinkers, we show that the blow-up limits of the nearby $\mathcal{M}_{s}$ move monotonically in the sense of rescaled flows (i.e. in a shrinker mean convex sense in the original scale). This built on previous work of the authors in [CM22], inspired by [ADS19, ADS20]. One then deduces from this, as in [Whi00], that the $(\mathcal{M}_{s})_{s\neq 0}$ only develop multiplicity-one spherical and cylindrical singularities and have genus zero near $(\mathbf{x}_{0},t_{0})$.
@@
1.3. The density-drop approach
As before, we let $\mathcal{M}_{s}$ be a monotone one-parameter family of flows, and we study blow-ups of this family at multiplicity-one singular points modeled by non-spherical, non-cylindrical shrinkers.
Let $\Sigma$ be any such shrinker. The key is Proposition 3.6, a qualitative geometric result characterizing such $\Sigma$: any translate of its shrinking spacetime track must “cross” its original, untranslated shrinking track. This is a special feature of shrinkers $\Sigma$ that are not spheres or cylinders. Clearly, it is not true for spheres and cylinders.
Our density-drop result, Proposition 4.1, quantifies Proposition 3.6. Let $\Theta_{\Sigma}$ be the Huisken density of the singular point modeled on $\Sigma$. Proposition 4.1 shows that every ancient flow to one side of the shrinking flow of $\Sigma$, with entropy no larger than $\Theta_{\Sigma}$, has Huisken density $\leq\Theta_{\Sigma}-\eta_{0}$ at all singular points (except the origin). Here, $\eta_{0}$ depends only on background a priori genus and entropy bounds, but not $\Sigma$. It follows from Proposition 4.1, and a monotonicity argument, that all sufficiently nearby points in the pre-blow-up picture have Huisken density $\leq\Theta_{\Sigma}-\eta_{0}$. Thus, our central singular point is isolated among singular points with density $\geq\Theta_{\Sigma}-\tfrac{1}{2}\eta_{0}$.
We now globalize. For each $\mathcal{M}_{s}$, $s=o(1)$, we may assume the existence of singular points that are not modeled by multiplicity-one spheres or cylinders. (Otherwise, we are done.) Let us then restrict our attention to the first (ordered by time) such singular point(s). Existence follows by work of Bernstein–Wang [BW17] (see Proposition 3.3). The work of Ilmanen [Ilm95b] and Bamler–Kleiner’s resolution of the multiplicity-one conjecture [BK23] (see Theorem 3.9) imply that these singular points are modeled by multiplicity-one shrinkers that are not spheres or cylinders. Therefore, Proposition 4.1 (density-drop) applies and guarantees that these points are isolated as long as we suitably stratify by density. In particular, by elementary considerations, the entire set of such points turns out to be finite, so we can clearly choose $\mathcal{M}_{s}$’s that miss it.
1.4. Comparison
The genus-drop approach to resolve Huisken’s conjecture is significantly longer than the density-drop in the current paper, but it also gives a refined local picture. In particular, together with the Bamler–Kleiner multiplicity-one theorem and their compactness theorem for bounded almost regular flows ([BK23, Theorem 1.7]), inner/outermost flows experience strict genus drop after every nongeneric singular point (thus bounding the number of nongeneric singular points by the initial genus) see, e.g., the proof of [BK23, Theorem 1.9 (c)].
@@ 1.6. Acknowledgements
OC was supported by a Terman Fellowship and an NSF grant (DMS-2304432). KC was supported by the KIAS Individual Grant MG078902. CM was supported by grant NSF DMS 2403728. We thank Richard Bamler and Bruce Kleiner for their interest in this work, and the anonymous referees for helpful suggestions.
@@ 2.1. Spacetime
Furthermore we denote parabolic distance in spacetime using
$d_{p}(X,Y)=d_{p}\big((\mathbf{x},t),(\mathbf{y},s)\big):=\sqrt{\|\mathbf{x}-\mathbf{y}\|^{2}+|t-s|},$
and the natural parabolic dilation around the spacetime origin $(\mathbf{0},0)$, by $\lambda>0$, using
$\operatorname{ParDil}_{\lambda}:\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R}^{3}\times\mathbb{R},\;\operatorname{ParDil}_{\lambda}(\mathbf{x},t)=(\lambda^{-1}\mathbf{x},\lambda^{-2}t).$
$2.$ $F$-functional and density
@@
which is nondecreasing in $r$ by Huisken’s well-known monotonicity formula [Hui90]. In particular,
@@ 2.3. Shrinkers
$\mathcal{S}$ $:=\{\Sigma\subset\mathbb{R}^{3}\text{ embedded shrinker with }\lambda(\Sigma)<\infty\},$
@@ 2.4. Brakke flows
    For a bounded interval $t_{1},t_{2}]\subset I$ and any compact set $K\subset\mathbb{R}^{3}$,
    $\int_{t_{1}}^{t_{2}}\int_{K}(1+|\mathbf{H}|^{2})d\mu(t)dt<\infty.$
@@
    If $t_{1},t_{2}]\subset I$, $f\in C^{1}_{c}(\mathbb{R}^{3}\times[t_{1},t_{2}])$, and $f\geq 0$, then Brakke’s inequality holds:
    $\int f(\cdot,t_{2})\,d\mu(t_{2})-\int f(\cdot,t_{1})\,d\mu(t_{1})\leq\int_{t_{1}}^{t_{2}}\int\left(-|\mathbf{H}|^{2}f+\mathbf{H}\cdot\nabla f+\tfrac{\partial}{\partial t}f\right)\,d\mu(t)\,dt.$
@@
In this weaker setting setting, (2.1), (2.2) extend to varifolds by looking at their induced Radon measures, and the monotonicity underlying (2.3), (2.4) extends to Brakke flows in $\mathbb{R}^{3}$ with bounded area ratios (cf. [Ilm95b, Lemma 7]).
Furthermore, we slightly abuse notation and still denote with $\operatorname{ParDil}_{\lambda}$ the standard parabolic dilation of a Brakke flow around the spacetime origin $(\mathbf{0},0)$ by a factor $\lambda>0$. This is obtained by pushing forward using the parabolic dilation map from Section 2.1.
@@ 2.5. $F$-stationary varifolds
The shrinker condition is also equivalent to the following (see [Ilm95a], [CM12a, §3]):
@@
A varifold $V$ in $\mathbb{R}^{3}$ is called $F$-stationary if is stationary with respect to the conformally flat metric $e^{-|\mathbf{x}|^{2}/4}g_{\mathbb{R}^{3}}$ on $\mathbb{R}^{3}$.
@@
$\tilde{\mathcal{M}}_{i}:=\operatorname{ParDil}_{\lambda_{i}}(\mathcal{M}-X)\rightharpoonup\tilde{\mathcal{M}}$,
@@
for some $F$-stationary integral varifold $V$ in $\mathbb{R}^{3}$. See [Ilm95b, Lemma 8] or [Whi97].
@@ 3.1. Shrinker compactness
$\operatorname{genus}(\Sigma_{i})\leq g,\;F(\Sigma_{i})\leq\Lambda,$
@@
$\Sigma_{i}\to\Sigma\text{ in }C^{\infty}_{\textnormal{loc}},\;\operatorname{genus}(\Sigma)\leq g,\;F(\Sigma_{i})\to F(\Sigma).$
@@ 3.4. Frankel property for $F$-stationary varifolds
then $V=k\Sigma$ for some $k\in\mathbb{N}$. [^3] Here and throughout, we implicitly identify $\Sigma$ with its induced integral varifold.
@@ 3.5. Geometric characterization of generic singularities
In short, Proposition 3.6 characterizes $\Sigma\in\mathcal{S}^{\textrm{gen}}$ as the only shrinkers having some nontrivial translate of their spacetime track (i.e., $X_{0}+\operatorname{supp}\mathcal{M}_{\Sigma}$ for some $X_{0}\neq(\mathbf{0},0)$) contained, for sufficienty negative times, on one side of the original shrinking spacetime track (i.e., $\cup_{t<0}\sqrt{-t}\bar{\Omega}$, which is a side of $\mathcal{M}_{\Sigma}$).
@@ 3.6. First nongeneric singular time
We recall the following notion from [CCMS24a]: [^4] In the reference, $T_{\textnormal{bad}}(\mathcal{M})$ is equivalently defined as the supremum of times $T$ such that all tangent flows at $X=(\mathbf{x},t)\in\operatorname{sing}\mathcal{M}$, $t<T$, are for negative times coincident with some $\mathcal{M}_{\Sigma}$, $\Sigma\in\mathcal{S}^{\textnormal{gen}}$.
@@ 3.7. Ilmanen and Bamler–Kleiner tangent flow structure theorem
The work of Ilmanen [Ilm95b] and Bamler–Kleiner [BK23] gives us a precise description of tangent flows at the first nongeneric time time:
@@
Let $\mathcal{M}$ be a cyclic unit-regular integral Brakke flow with $\mathcal{M}(0)=\mathcal{H}^{2}\lfloor M$, where $M\subset\mathbb{R}^{3}$ is a closed embedded surface. If $T_{\textnormal{bad}}(\mathcal{M})<\infty$ and $X$ is a first nongeneric point for $\mathcal{M}$ (see Corollary 3.8), then every tangent flow to $\mathcal{M}$ at $X$ coincides, for $t<0$, with $\mathcal{M}_{\Sigma}$ for some $\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}$.
@@
The weaker result, with $\mathcal{M}_{\Sigma}$ replaced by $\mathcal{M}_{\Sigma}, kk\in\mathbb{N}$, and $\Sigma\in\mathcal{S}$, was proven in [Ilm95b] for the first singular time. Modifications of the proof to the first nongeneric singular time were give in the proof of [CCMS24a, Proposition 11.3].
The fact that $k=1$ and $\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}$ follows from Bamler–Kleiner’s recent resolution of the multiplicity-one conjecture. In particular, it follows from [BK23, Theorem 1.2] provided we can ensure $\mathcal{M}$ is what Bamler–Kleiner call “almost-regular” on $0,T_{\textnormal{bad}}(\mathcal{M}))$. This is justified in [BK23, Lemma 7.8]. ∎
@@
This is our main density-drop result. It quantifies the characterization in Proposition 3.6. Subject to appropriate genus and entropy bounds, ancient cyclic unit-regular integral Brakke flows on one side of a spacetime track of a non-generic shrinker have quantitatively lower densities relative to the density of the non-generic shrinker.
Proof of Proposition 4.1.
@@
Using the compactness of cyclic unit-regular integral Brakke flows ([Ilm94] for integral Brakke flows, [Whi05] for unit-regularity, and [Whi09] for cyclicity) together with (c) and (4.4), we can pass (c) to the limit:
@@
The upper semicontinuity of densities together with (4.1), (4.4), (4.6), (4.7), (4.8), implies
@@
In particular, $\mathcal{M}$ is a shrinking flow with spacetime center $X$ due to the monotonicity formula. If $V$ is the $F$-stationary varifold corresponding to some tangent flow to $\mathcal{M}$ at -$\infty$, then
@@
It follows from Proposition 3.5, (4.10), and (4.11), that $V=\Sigma$. This forces Proposition 3.6 into a contradiction with the fact that $\Sigma\not\in\mathcal{S}^{\textnormal{gen}}$ per (4.5) and $X\neq(\mathbf{0},0)$ (since $d_{p}(X,(\mathbf{0},0))=1$) per (4.6). This completes our proof. ∎
@@
For each $s\in(-1,1)$, denote $M_{s}:=\partial K_{s}$ and let $\mathcal{M}_{s}$ be any cyclic unit-regular integral Brakke flow with $\mathcal{M}_{s}(0)=\mathcal{H}^{2}\lfloor M_{s}$ (see [Ilm94]). [^5] In our previous work, we have often had to consider multiple Brakke flows for each fixed $s\in(-1,1)$. Our more streamlined Claim 4.2 allows us to consider just a single Brakke flow for our argument.
@@
Claim 4.2.
Fix $s_{0}\in(-1,1)$ and $X_{0}$ a first nongeneric point for $\mathcal{M}_{s_{0}}$ (recall Corollary 3.8). If $s_{i}\to s_{0}$, and $X_{0}\neq X_{i}\to X_{0}$ are first nongeneric points for $\mathcal{M}_{s_{i}}$, then
@@
for sufficiently large $i$. Translate so that $X_{0}=(\mathbf{0},0)$ and set $\lambda_{i}:=d_{p}(X_{i},(\mathbf{0},0))>0$.
@@
(4.15) $\tilde{\mathcal{M}}_{s_{0}}^{i}:=\operatorname{ParDil}_{\lambda_{i}}\mathcal{M}_{s_{0}}\rightharpoonup\tilde{\mathcal{M}}_{s_{0}}=\mathcal{M}_{\Sigma}\text{ for some }\Sigma\in\mathcal{S}^{*}\setminus\mathcal{S}^{\textnormal{gen}}.$
@@
By the nonfattening of the flow before $X_{0}$ by Hershkovits–White and Choi–Haslhofer–Hershkovits [HW20, CHH22], the regularity for a.e. $t$ of flows with only generic singularities by [Whi97] (see also [CM16]), and White’s genus monotonicity [Whi95],
@@
$\tilde{\mathcal{M}}^{i}:=\operatorname{ParDil}_{\lambda_{i}}\mathcal{M}_{s_{i}}\rightharpoonup\tilde{\mathcal{M}}$,
@@
$=\lim_{i}\Theta_{\mathcal{M}_{i}}(\operatorname{ParDil}_{\lambda_{i}}Y,\lambda_{i}r)$
$\leq\lim_{i}\Theta_{\mathcal{M}_{i}}(\operatorname{ParDil}_{\lambda_{i}}Y,\rho)$
@@
where $\Omega\subset\mathbb{R}^{3}$ is a component of $\mathbb{R}^{3}\setminus\Sigma$. This follows from (4.12) combined with the avoidance principle and the locally smooth convergence $\tilde{\mathcal{M}}^{i}_{s_{0}}\rightharpoonup\mathcal{M}_{\Sigma}$ on $\mathfrak{t}^{-1}((-\infty,0))$ from (4.15) and Brakke’s regularity theorem [Bra75], the latter of which applies since $\Sigma$ is smooth.
Altogether (4.16), (4.17), (4.18), (4.19), (4.20), and $\tilde{X}\neq(\mathbf{0},0)$ (since $d_{p}(\tilde{X},(\mathbf{0},0))=1$) contradict our choice of $\eta_{0}$ coming from Proposition 4.1. This completes our proof of the claim. ∎
@@
 $\qquad\text{ and }\Theta_{\mathcal{M}_{s}}(X)\in[1+\tfrac{1}{2}\ell\eta_{0},1+\tfrac{1}{2}(\ell+1)\eta_{0})\}.$
The claim implies that $\mathcal{S}_{\ell}$ is discrete. In particular, $\mathcal{S}_{\ell}$ is finite since the flows become extinct after a uniform finite time. Therefore, $\cup_{\ell}\mathcal{S}_{\ell}$ is also finite since the union can be taken over a finite set of $\ell$’s by (4.13). Thus, we have that:
[…]
is finite.